OpenStudy (anonymous):

(2x10 to 3rd power) x (3x10 to eh 5th power)

OpenStudy (anonymous):

do you mean \[(2 \times 10^{3}) \times (3 \times 10^{5})\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If so you need to remember the operations with exponents rule?

OpenStudy (anonymous):

I thought so, but my question said without exponents.

OpenStudy (anonymous):

okay mulipty the numbers 2 times 3 to get 6

OpenStudy (anonymous):

okay then how would you solve this \[10^{3} \times 10^{5}\]

OpenStudy (anonymous):

add the exponents! to get \[10^{8}\]

OpenStudy (anonymous):

check is you 2 times 3 is between 1 and 9, and it is so your answer is \[6 \times 10^{8}\]

OpenStudy (anonymous):

I am sorry my computer froze on me.

OpenStudy (anonymous):

okay did you see the nswer

OpenStudy (anonymous):

Yes. I understood the 2 times 3 = 6 However, how did you get 108

OpenStudy (anonymous):

Okay I understand how you reached 10 by 8the power.

OpenStudy (anonymous):

I add ther exponents, the rule for multiplying exponents with the same base is to add the exponents

OpenStudy (anonymous):

great